OpenStudy (anonymous):

A car is traveling at a rate of 25 meters per second. What is the car's rate in kilometers per hour? How many kilometers will the car travel in 3 hours? Do not round your answers.

OpenStudy (anonymous):

\[\frac{ 25 meters }{ 1 second } \times \frac{ 60 seconds }{ 1 minute } \times \frac{ 60 minutes }{ 1 hour } \times \frac{ 1 kilometer }{ 1000 meters }\] every unit ancels out exept for kilimoter and hour

OpenStudy (anonymous):

my c button is iffy lol. cancels* except*